# Li He1, Pei-Yang Lin1, Peng-Hu Wei5, Guo-Guang Zhao5, Yun-Zhe Liu†,1,4 1

### v1

@@ 4.2 Implementation Details
The length of the sEEG patch is 100ms, resulting in 40 patches per sample in the pre-training dataset and 30 patches per sample in the downstream dataset. The "Spatial Encoder" contains one linear projection and three 1-D convolution layers, transforming the original sEEG patches into patch embeddings with $d=160$. The following "Transformer Encoder" contains an 8-layer Transformer encoder with model dimension $d=160$, inner dimension (FFN) $d_{ff}=320$, and 8 attention heads. See Appendix C for more details.
@@
During the pre-training, we use either all sEEG recordings (15 hours) or the sEEG recordings without task recordings (12 hours) to train the Du-IN VQ-VAE and Du-IN MAE models. To enhance the robustness of learned codexes and representations, we further use data augmentation described in Appendix D. For each subject, the model is pre-trained on a Linux system with 2 CPUs (Intel Xeon Gold 6230 40-Core Processor) and 1 GPU (NVIDIA Tesla V100 32GB) for $\sim$ 1.2 days.
@@
During the downstream evaluation, we split the task recordings into training, validation, and testing splits with a size roughly proportional to 80%, 10%, and 10%. We also use data augmentation, as described in Appendix D, to make the most of the gathered dataset. We employ cross-entropy loss (multi-class classification) as the training loss. Our experiments are conducted on one V100 GPU by Python 3.11.7 and PyTorch 2.1.2 + CUDA 12.3. The best models are trained based on the training set, selected from the validation set, and finally evaluated on the test set. For model comparison, we report the average and standard error values (of all subjects) on six different random seeds to obtain comparable results. For the results of the subject-wise evaluation, we report the average and standard deviation values (of each subject) in Appendix I.
@@ 4.3 Channel Contribution and Selection
To further understand the detailed contribution of each channel, we analyze the weights of linear projection in the spatial encoder. In detail, we calculate the contribution scores of channels per subject and organize them accordingly, as described in Appendix H. Figure 4 demonstrates that (1) the brain regions effective for speech decoding align with findings from previous neuroscience research, and (2) our model achieves optimal decoding performance with approximately 10 channels, 80% of which originate from the same electrode. To streamline, we utilize these top 10 channels for both pre-training and downstream evaluation.
@@ 4.4 Comparasion with Other Models
Table 2 presents the results of our Du-IN model, previous sEEG self-supervised models, and other supervised baselines. The results demonstrate that our Du-IN model outperforms all baselines. It’s worth noting that the models (i.e., BrainBERT, Brant) that build univariate representations based on a single channel perform worse than the models building multi-variate representations, challenging the generalizability of current strategies to model spatial relationships among channels with Transformer.
As the BrainBERT model doesn’t consider the spatial relationships among channels, we mainly focus on understanding why Brant fails to capture the spatial relationships on the speech decoding task effectively. Compared to epilepsy detection, speech is a more difficult task, demanding intricate processing in specific brain regions. However, unlike LaBraM jiang2024large , Brant doesn’t introduce spatial embeddings to identify the spatial location of each channel. Since the electrodes are sparsely distributed in the brain and the raw sEEG signals on the same electrode are highly correlated, it’s fairly easy to identify their spatial relationships through their values. As demonstrated in iTransformer liu2023itransformer , this modeling approach is well suited for detecting time-delay events, e.g., epilepsy detection.
For speech decoding tasks, sEEG often requires bi-polar re-reference (or Laplacian re-reference) to remove the high correlations among channels, thus avoiding model overfitting wang2023brainbert . Once the correlations among channels have been removed, Brant will lose the ability to model spatial relationships among channels. Meanwhile, Brant only involves a few sEEG location configurations (i.e., 10 subjects) during its pre-training stage. Based on their methods, more sEEG location configurations should be required to learn generic relationships among channels.
@@
Methods Pre-trained Model Size Accuracy (%) $\pm$ Ste (%)
EEG-Conformersong2022eeg 2.34M 45.82$\pm$4.66
CNN-BiGRUmoses2021neuroprosthesis 0.54M 32.04$\pm$5.45
BrainBERTwang2023brainbert 43.58M 7.50$\pm$1.76
Brantzhang2024brant 69.35M 12.42$\pm$4.10
Du-IN 4.38M 56.29$\pm$5.20
Du-IN (vqvae+vq) 4.38M 44.17$\pm$4.04
Du-IN (vqvae) 4.38M 58.24$\pm$4.83
Du-IN (mae) 4.38M 62.70$\pm$4.69
@@ 4.5 Ablation Study
Table 3: Ablation study on whether pre-training with the downstream dataset or not.
@@
Du-IN (mae w/o downstream dataset) 12 hours per subject 60.02$\pm$4.34
Du-IN (mae w/ downstream dataset) 15 hours per subject 62.70$\pm$4.69
@@
During the neural tokenizer training stage, the Du-IN VQ-VAE model encodes sEEG patches into discrete codexes and then reconstructs the original signal from these codexes. We evaluate performance against varying codex sizes (512 to 8192) to ascertain if codex size affects the quality of the learned codebook. As illustrated in Figure 5, while extremely small codex size lacks representation diversity, extremely large codex size often leads to codebook collapse. We suspect that our existing training data might not be adequate for larger codex sizes. Furthermore, our experiments suggest that the model performs optimally when the codex dimension, denoted as $d_{codex}=64$, is slightly less than the model dimension, $d=160$, yielding a more effective regularization effect.
@@
We also conduct the ablation study on the model structure for the spatial encoder described in Section 3.2. As the spatial encoder transforms the sEEG signals within a given patch to a patch embedding, it compresses the sEEG signals for perception. As described in Section 4.2, the model utilizes a perception field of 100ms. We conduct an ablation study of different perception fields and report it in Figure 5. The model performance notably drops with a perception field smaller than 60ms and gradually declines as the perception field exceeds 160ms. The model reaches a small peak around 100ms to 140ms. We think this phenomenon is rational since sEEG is known for its ability to capture the rapid dynamics of specific brain regions precisely.
Figure 5: Ablation study on different codex sizes, codex dimensions, and perception fields.
@@ 5 Limitations
Despite Du-IN’s enhancements in speech decoding via discrete codebook-guided mask modeling, it is still restricted to close-set speech decoding tasks (i.e., the word set only includes 61 pre-determined words). However, a parallel to our work, Feng et al. feng2023high build an acoustic-inspired model that can decode arbitrary Chinese words by predicting syllable components (initials, finals, tones). Although their method requires a large amount of labeled data, their experimental design mirrors ours closely. The difference lies in the requirement for the subject to repeat syllable components, instead of entire words. Therefore, with slight modifications, our model can support open-set speech decoding tasks.
Additionally, the experiments in this paper are restricted to the vocal production part of language decoding, i.e. speech decoding. A more interesting but difficult task is to decode language from the semantic level, in which large language models have been wildly used to improve the model performance tang2023semantic ; duan2023dewave . However, due to the locality of sEEG recordings, it is still under exploration whether sEEG recordings can fully capture semantic-related information across brain regions.
@@ 6 Conclusion
This paper proposes Du-IN, a framework for speech decoding, which learns contextual embeddings through discrete codebook-guided mask modeling on specific brain regions. To evaluate our model, we collect a well-annotated Chinese word-reading sEEG dataset to address the lack of sEEG language dataset. Inspired by neuroscientific findings, we analyze the effective brain regions for speech decoding and achieve the best decoding performance with about one electrode in specific brain regions, which dovetails with the past neuroscientific research on language. Comprehensive experiments demonstrate that our model outperforms both supervised and sEEG-based self-supervised baselines, effectively capturing the intricate processing within specific brain regions. It marks a promising neuro-inspired AI approach in BCI. In the end, we hope our work can have implications for future developments in sEEG-based self-supervised models with more consideration over how to build the basic representation units so that the model can maximally benefit from the pre-training stage.
@@
Reproducibility Statement
Code to train models and reproduce the results is submitted as part of the supplementary materials and can be accessed here: TODO, including a demo dataset of 3 subjects for downstream fine-tuning.
@@ Appendix B Details of Baselines
  • EEG-Conformer[34]: A supervised model that consists of both CNN module and Transformer module, to encapsulate local and global features in a unified EEG classification framework. EEG-Conformer is mainly designed for EEG-based motor imagination tasks. Since the data modes of EEG and sEEG are similar, and the signals primarily pertain to vocal production, EEG-Conformer is suitable to serve as a baseline for comparison.
@@
  • BrainBERT[38]: A self-supervised model for sEEG recordings that bridges modern representation learning approaches to neuroscience. BrainBERT builds universal representation based on the superlet spectrograms of one single sEEG channel without modeling the spatial relationships among channels. Since the downstream tasks for BrainBERT are also related to language decoding (e.g., sentence-onset detection, speech vs. non-speech detection, etc.), BrainBERT is suitable to serve as a baseline for comparison.
@@
  • Brant[41]: A self-supervised model for sEEG recordings that can capture both long-term temporal dependency and spatial correlation from neural signals. Brant is mainly designed for medicine, serving as a sEEG foundation model. Although Brant mainly evaluates its performance on the low-level modeling tasks [39] (e.g., neural signal forecasting, imputation, etc.), Brant achieves SOTA performance on some high-level modeling tasks (e.g., seizure detection). As a foundation model in sEEG pre-training field, Brant is suitable to serve as a baseline for comparison.
@@
  • For the BrainBERT model, we use around 180 hours of sEEG recordings from all 12 subjects for pre-training. This pre-training dataset is larger than the one (approximately 45 hours) used in the original paper.
  • For the Brant model, we also use all sEEG recordings from 12 subjects to pre-train it. While the total pre-training dataset is smaller than the one (around 2700 hours) used in the original paper, the number of subjects (i.e., the number of sEEG location configurations) is greater than in the original paper.

### v3

@@ 4.2 Implementation Details
The length of the sEEG patch is 100ms, resulting in 40 patches per sample in the pre-training dataset and 30 patches per sample in the downstream dataset. The "Spatial Encoder" contains one linear projection and three 1D convolution layers, transforming the original sEEG patches into patch embeddings with $d=160$. The following "Transformer Encoder" contains an 8-layer Transformer encoder with model dimension $d=160$, inner dimension (FFN) $d_{ff}=320$, and 8 attention heads. See Appendix C for more details.
@@
During the pre-training, we use either all sEEG recordings (15 hours) or the sEEG recordings without task recordings (12 hours) to train the Du-IN VQ-VAE and Du-IN MAE models. To enhance the robustness of the learned codex and representations, we further use data augmentation described in Appendix D. For each subject, the model is pre-trained on a Linux system with 2 CPUs (Intel Xeon Gold 6230 40-Core Processor) and 1 GPU (NVIDIA Tesla V100 32GB) for $\sim$ 1.2 days.
@@
During the downstream evaluation, we split the task recordings into training, validation, and testing splits with a size roughly proportional to 80%, 10%, and 10%. All experiments are conducted on the same machine with the same set of random seeds. The train/validation/test splits are the same across different models. We also use data augmentation, as described in Appendix D, to make the most of the gathered dataset. We employ cross-entropy loss (multi-class classification) as the training loss. Our experiments are conducted on one V100 GPU by Python 3.11.7 and PyTorch 2.1.2 + CUDA 12.3. The best models are trained based on the training set, selected from the validation set according to accuracy, and finally evaluated on the test set. For model comparison, we report the average and standard error values (of all subjects) on six different random seeds to obtain comparable results. For the results of the subject-wise evaluation, we report the average and standard deviation values (of each subject) in Appendix K.
@@ 4.3 Channel Contribution and Selection
To further understand the detailed contribution of each channel, we analyze the weights of linear projection in the spatial encoder. In detail, we calculate the contribution scores of channels per subject and organize them accordingly, as described in Appendix H. Figure 4 demonstrates that (1) the brain regions effective for speech decoding align with findings from previous neuroscience research, and (2) our model achieves optimal decoding performance with approximately 10 channels, 80% of which originate from the same electrode. To streamline, we utilize these top 10 channels (selected according to train-set) for both pre-training and downstream evaluation.
@@ 4.4 Comparasion with Other Models
Table 2 presents the results of our Du-IN model and the advanced baselines that are designed for either brain signals or general time series. See Appendix B and Appendix C.3 for detailed descriptions of models. The results demonstrate that our Du-IN model outperforms all baselines. It’s worth noting that the models (i.e., the foundation models designed for brain signals) that adopt spatial-temporal integration to model spatial relationships among channel-level tokens perform worse than the models that adopt temporal modeling based on region-level tokens, challenging the generalizability of current strategies to model spatial relationships among channels with Transformer.
@@
Methods Token Level Config Model Size Accuracy (%) $\pm$ Ste (%)
PT11footnotemark: 1 MS22footnotemark: 2
TS-TCCeldele2021time Region 0.32M 24.85$\pm$4.42
CNN-BiGRUmoses2021neuroprosthesis Region - 0.54M 32.04$\pm$5.45
EEG-Conformersong2022eeg Region - 2.34M 45.82$\pm$4.66
Neuro-BERTwu2024neuro Region 2.14M 49.51$\pm$4.43
DeWaveduan2023dewave Region - 5.70M 32.43$\pm$4.48
BrainBERTwang2023brainbert Channel 43.58M 6.72$\pm$1.59
BrainBERTwang2023brainbert Channel 43.58M 7.50$\pm$1.76
Brantzhang2024brant Channel 69.35M 11.16$\pm$3.56
Brantzhang2024brant Channel 69.35M 12.42$\pm$4.10
LaBraMjiang2024large Channel 6.85M 11.53$\pm$2.63
LaBraM-PopTjiang2024large ; chau2024population Channel 6.85M 11.78$\pm$2.70
Du-IN Region - 4.38M 56.29$\pm$5.20
Du-IN (vqvae+vq) Region 4.38M 44.17$\pm$4.04
Du-IN (vqvae) Region 4.38M 58.24$\pm$4.83
Du-IN (mae) Region 4.38M 62.70$\pm$4.69
Du-IN (poms) Region 5.18M 59.18$\pm$4.63
  • 1
    PT: Whether the model is pre-trained before evaluation.
  • 2
    MS: Whether the model is pre-trained across multiple subjects.
As BrainBERT wang2023brainbert doesn’t consider the spatial relationships among channels, we mainly focus on understanding why Brant zhang2024brant , LaBraM jiang2024large and LaBraM-Popt jiang2024large ; chau2024population fail to effectively capture the discriminative features on the speech decoding task. These models typically build channel-level tokens by segmenting non-overlapping patches with large receptive fields (e.g., 1 second) from single channels. However, this approach makes it challenging to capture the rapid process of brain dynamics. Moreover, while these models further utilize Transformer to capture the spatial relationships among these tokens, they do not encourage region-level embeddings, either through their architecture yi2024learning or their pre-training objective chau2024population . Therefore, the effectiveness of building brain foundation models based on these spatial-temporal backbones is still under exploration, especially for cognitive tasks (e.g., speech decoding), which are of great value in the field of neuroscience.
Besides, unlike LaBraM jiang2024large , Brant doesn’t introduce spatial embeddings to identify the spatial location of each channel. Since the electrodes are sparsely distributed in the brain and the raw sEEG signals on the same electrode are highly correlated, it’s fairly easy to identify their spatial relationships through their values. As demonstrated in iTransformer liu2023itransformer , this modeling approach is well suited for detecting time-delay events, e.g., seizure detection. For speech decoding tasks, sEEG often requires bi-polar re-reference (or Laplacian re-reference) to remove the high correlations among channels, thus avoiding model overfitting wang2023brainbert . Once the correlations among channels have been removed, Brant will lose the ability to model spatial relationships among channels.
For other baselines that use temporal modeling based on region-level tokens, we provide a detailed explanation of their performance differences as follows. TS-TCC eldele2021time tokenizes raw sEEG signals into region-level tokens with a stack of 1D depthwise convolution blocks, but it lacks a temporal Transformer for further integration over time. CNN-BiGRU moses2021neuroprosthesis introduces a stack of GRU layers on top of these tokens to perform temporal integration. EEG-Conformer song2022eeg introduces a temporal Transformer to better integrate global temporal information, which makes it outperform CNN-BiGRU. However, EEG-Conformer tokenizes raw sEEG signals with the temporal-spatial convolution, applying the same convolutional kernel across different channels, which overlooks the specificity of brain computation caucheteux2023evidence . This also raises a challenge for the effectiveness of current sEEG foundation models, which rely on shared convolution blocks across individual channels. Neuro-BERT wu2024neuro further introduces mask modeling to learn contextual embeddings, which makes it outperform EEG-Conformer. DeWave duan2023dewave utilizes the Conformer model gulati2020conformer for tokenization, which involves more parameters but is less effective than 1D depthwise convolution.
@@ 4.5 Ablation Study
Table 3: Ablation study on whether pre-training with the downstream dataset (DD) or not.
@@
Du-IN (mae w/o DD) 12 hours per subject 60.02$\pm$4.34
Du-IN (mae w/ DD) 15 hours per subject 62.70$\pm$4.69
@@
During the Du-IN VQ-VAE training stage, the Du-IN VQ-VAE model encodes sEEG patches into discrete codes and then reconstructs the original signal from these codes. We evaluate performance against varying codex sizes (512 to 8192) to ascertain if codex size affects the quality of the learned codex. As illustrated in Figure 5, while extremely small codex size lacks representation diversity, extremely large codex size often leads to codex collapse. We suspect that our existing training data might not be adequate for larger codex sizes. Furthermore, our experiments suggest that the model performs optimally when the codex dimension, denoted as $d_{codex}=64$, is slightly less than the model dimension, $d=160$, yielding a more effective regularization effect.
@@
We also conduct the ablation study on the model structure for the spatial encoder described in Section 3.2. As the spatial encoder transforms the sEEG signals within a given patch to a patch embedding, it compresses the sEEG signals for perception. As described in Section 4.2, the model utilizes a receptive field of 100ms. We conduct an ablation study of different receptive fields and report it in Figure 5. The model performance notably drops with a receptive field smaller than 60ms and gradually declines as the receptive field exceeds 160ms. The model reaches a small peak around 100ms to 140ms. We think this phenomenon is rational since sEEG is known for its ability to capture the rapid dynamics of specific brain regions precisely.
Figure 5: Ablation study on different codex sizes, codex dimensions, and receptive fields.
@@ 5 Limitations
Despite Du-IN’s enhancements in speech decoding via discrete codex-guided mask modeling, it is still restricted to close-set speech decoding tasks (i.e., the word set only includes 61 pre-determined words). However, a parallel to our work, Feng et al. feng2023high , which follows previous works herff2015brain ; soroush2023nested , build an acoustic-inspired model that can decode arbitrary Chinese words by predicting syllable components (initials, finals, tones). Although their method requires a large amount of labeled data, their experimental design mirrors ours closely. The difference lies in the requirement for the subject to repeat syllable components, instead of entire words. Therefore, with slight modifications, our model can support open-set speech decoding tasks.
Additionally, the experiments in this paper are restricted to the vocal production part of language decoding, i.e., speech decoding. A more interesting but difficult task is to decode language from the semantic level, in which large language models have been wildly used to improve the model performance tang2023semantic ; duan2023dewave . However, due to the locality of sEEG recordings, it is still under exploration whether sEEG recordings can fully capture semantic-related information across brain regions.
@@ 6 Conclusion
This paper proposes Du-IN, a framework for speech decoding, which learns contextual embeddings through discrete codex-guided mask modeling on specific brain regions. To evaluate our model, we collect a well-annotated Chinese word-reading sEEG dataset to address the lack of sEEG language dataset. Inspired by neuroscientific findings, we analyze the effective brain regions for speech decoding and achieve the best decoding performance with about one electrode in specific brain regions, which dovetails with the past neuroscientific research on language. Comprehensive experiments demonstrate that our model outperforms both supervised and sEEG-based self-supervised baselines, effectively capturing the intricate processing within specific brain regions. It marks a promising neuro-inspired AI approach in BCI. In the end, we hope our work can have implications for future developments in sEEG-based self-supervised models with more consideration over how to build the basic representation units so that the model can maximally benefit from the pre-training stage.
7 Broader Impacts
Our method advances the feasibility of invasive BCI technology by being the first to demonstrate speech decoding using a single sEEG electrode, which holds significant potential for clinical applications. For patients who have lost their ability to communicate or perform daily tasks due to neurological conditions like locked-in syndrome or amyotrophic lateral sclerosis (ALS), our approach offers a less invasive alternative to technologies like ECoG or microelectrode arrays, thereby reducing the risk of brain damage.
Acknowledgements
This study was supported by the National Science and Technology Innovation 2030 Major Program (2022ZD0205500), the National Natural Science Foundation of China (32271093), the Beijing Natural Science Foundation (Z230010, L222033), and the Fundamental Research Funds for the Central Universities. We would like to extend our sincere appreciation to Dr. Zhi-Feng Yue for his coordination and support in securing the computing resources essential for this study. Besides, we also sincerely appreciate the LaBram jiang2024large team for their valuable discussion on the visual design of Figure 2 and Figure 3.
@@ Acknowledgements
Our informed consent signing process is as follows:
  1. 1.
    If the experimental participants are adults and have full civil capacity, we will ask them to sign a written informed consent after the participants have fully informed consent;
  2. 2.
    If the experimental participants are minors or do not have full civil capacity, we will ask the participant’s legal guardian to sign a written informed consent after the participants and their legal guardians have fully informed consent.
Our informed consent form includes the following points:
  1. 1.
    Contact information of research institutions and researchers;
  2. 2.
    Research direction and purpose;
  3. 3.
    Risks involved in the research;
  4. 4.
    Personal information, data and usage methods to be used in the research;
  5. 5.
    Privacy protection statement (all personal identification information (PII) will not be disclosed);
  6. 6.
    Data storage statement (retained after deleting all personal identification information (PII));
  7. 7.
    Voluntary statement of participants;
  8. 8.
    Statement that participants can withdraw unconditionally at any time.
Our data storage and protection procedures include the following processes:
  1. 1.
    Our data collection, transfer, and analysis tasks are only completed by researchers who have signed relevant confidentiality agreements;
  2. 2.
    The collected raw data will be copied twice as soon as possible, one copy to a storage computer that is not connected to the Internet and encrypted, and the other copy to a mobile hard disk and encrypted and stored offline;
  3. 3.
    The use of the data is only authorized to the research leader and the main researchers (less than 5 people), among which the main researchers can only access data that does not contain personal identification information (PII);
  4. 4.
    After the study is completed, all personal identification information (PII) on both nodes (storage computer, mobile hard disk) will be deleted immediately.
To prevent unauthorized access or possible data leakage, we use double encryption on the storage computer, that is, a static password and a dynamic password (received by mobile phone or email); physical isolation is used on the mobile hard disk, that is, it is locked in a filing cabinet, and the key is only kept by the research leader and the main researchers.
@@ Appendix B Details of Baselines
  • TS-TCC[20]: A self-supervised model that consists only of a CNN module to capture local features. This model learns robust temporal and discriminative representations from time series by designing a tough cross-view prediction task and a contextual contrasting module. Since sEEG is a unique type of time series, this model is suitable to serve as a baseline for comparison.
@@
  • EEG-Conformer[44]: A supervised model that consists of both CNN module and Transformer module, to encapsulate local and global features in a unified EEG classification framework. EEG-Conformer is mainly designed for EEG-based motor imagination tasks. Since the data modes of EEG and sEEG are similar, and the signals primarily pertain to vocal production, this model is suitable to serve as a baseline for comparison.
  • Neuro-BERT[50]: A self-supervised model that consists of both CNN module and Transformer module, to encapsulate local and global features. This model learns robust contextual representations from EEG by introducing mask modeling. Since the data modes of EEG and sEEG are similar, this model is suitable to serve as a baseline for comparison.
  • DeWave[19]: A supervised model that consists of both Conformer module [24] and Transformer module, to encapsulate local and global features for language decoding. We adopt its encoder, which consists of a 6-layer Conformer and a 6-layer Transformer. Then, we add a classification head, which is also used in our model, for downstream word classification. Since DeWave is also designed for language decoding, this model is suitable to serve as a baseline for comparison.
@@
  • BrainBERT[49]: A self-supervised model for sEEG recordings that bridges modern representation learning approaches to neuroscience. BrainBERT builds universal representation based on the superlet spectrograms of one single sEEG channel without modeling the spatial relationships among channels. Since the downstream tasks for BrainBERT are also related to language decoding (e.g., sentence-onset detection, speech vs. non-speech detection, etc.), this model is suitable to serve as a baseline for comparison.
@@
  • Brant[54]: A self-supervised model for sEEG recordings that can capture both long-term temporal dependency and spatial correlation from neural signals. Brant is mainly designed for medicine, serving as a sEEG foundation model. Although Brant mainly evaluates its performance on the low-level modeling tasks [51] (e.g., neural signal forecasting, imputation, etc.), Brant achieves SOTA performance on some high-level modeling tasks (e.g., seizure detection). As a foundation model in sEEG pre-training field, this model is suitable to serve as a baseline for comparison.
  • LaBraM[28]: A self-supervised model for EEG recordings that learns generic representations with tremendous EEG data. LaBraM serves as an EEG foundation model, achieving SOTA performance on various downstream EEG tasks. Since the spatial embeddings are pre-defined according to the EEG caps, LaBraM can only be trained within one subject under the sEEG setting. Since the data modes of EEG and sEEG are similar, this model is suitable to serve as a baseline for comparison.
  • LaBraM+PopT[28, 10]: A self-supervised model based on LaBraM, simply replacing the learnable spatial embeddings with hard-coded spatial embeddings from PopT [10] to enable multi-subject pre-training under the sEEG setting.
The detailed implementations of these baseline models are given here:
  • For the TS-TCC method [20], the hyper-parameters are optimized for better performance, as they also have different hyper-parameter settings for different datasets in their original implementation. The data samples are resampled to 400Hz. The sizes of convolution kernels are changed to {25, 8, 8} (other attempts include {8, 8, 8}, {15, 8, 8}, {20, 8, 8}, and {30, 8, 8}); the sizes of pooling kernels are changed to {10, 2, 2} (other attempts include {2, 2, 2}, {5, 2, 2}, and {20, 2, 2}); the numbers of pooling strides are changed to {10, 2, 2} (other attempts include {2, 2, 2}, {5, 2, 2}, and {20, 2, 2}). All other hyper-parameters are the same as the original implementation.
  • For the CNN-BiGRU method [37], the hyper-parameters are the same as the original implementation. The data samples are resampled to the specified sampling rate.
  • For the EEG-Conformer method [44], the hyper-parameters are the same as the original implementation. The data samples are resampled to the specified sampling rate.
  • For the Neuro-BERT method [50], the hyper-parameters are optimized for better performance, as they also have different hyper-parameter settings for different datasets in their original implementation. The data samples are sampled to 400Hz. The sizes of convolution kernels are changed to {40,} (other attempts include {20,} and {80,}); the numbers of convolution strides are changed to {40,} (other attempts include {20,} and {80,}).
  • For the DeWave method [19], the hyper-parameters are the same as the original implementation. The data samples are resampled to the specified sampling rate.
  • For the BrainBERT method [49], the hyper-parameters are optimized for better performance. We change the "nperseg" and "noverlap" arguments of "scipy.signal.stft" function from {400, 350} to {1600, 1400} (other attempts include {200, 175}, {800, 700} and {3200, 2800}).
  • For the Brant method [54], the hyper-parameters are optimized based on the Brant-Tiny model for better performance. We change the length of the patch segment from 6 seconds to 1 second. Besides, we change the linear embedding layer to the convolution embedding layer, which is also used in LaBraM [28]. The numbers of convolution filters are {96, 96, 96} (other attempts include {192, 192, 192}); the sizes of convolution kernels are {9, 9, 3} (other attempts include {19, 9, 3} and {9, 9, 3}); the numbers of convolution strides are {5, 5, 1} (other attempts include {10, 5, 1}) and {5, 5, 2}).
  • For the LaBraM method [28], the hyper-parameters are the same as the original implementation of the LaBraM-Base model. The data samples are resampled to the specified sampling rate.
  • For the LaBraM-PopT method [28, 10], the hyper-parameters are the same as the original implementation of the LaBraM-Base model. The data samples are resampled to the specified sampling rate.
@@
  • For the TS-TCC model, we use all sEGG recordings for each subject to pre-train it. The data samples are 4 seconds.
  • For the Neuro-BERT model, we use all sEGG recordings for each subject to pre-train it. The data samples are 4 seconds.
  • For the BrainBERT model, we use around 180 hours of sEEG recordings from either each subject or 12 subjects for pre-training. This pre-training dataset is larger than the one (approximately 45 hours) used in the original paper. The data samples are 4 seconds.
  • For the Brant model, we also use all sEEG recordings from either each subject or 12 subjects to pre-train it. While the total pre-training dataset is smaller than the one (around 2700 hours) used in the original paper, the number of subjects (i.e., the number of sEEG location configurations) is greater than in the original paper. The data samples are 4 seconds.
  • For the LaBraM model, we use all sEGG recordings for each subject to pre-train it. The data samples are 4 seconds.
  • For the LaBraM-PopT model, we use all sEEG recordings from 12 subjects to pre-train it. The data samples are 4 seconds.